I need to find the intersection points of the two equations x(squared)+y(squared)=5 and x-y=1. Can you help me with it?
Answer by stanbon(75887)

I need to find the intersection points of the two equations x(squared)+y(squared)=5 and x-y=1.
------------------------------
x^2+y^2=5
x=y+1
Substitute to get:
(y+1)^2+y^2=5
2y^2+2y-4=0
y^2+y-2=0
(y+2)(y-1)=0
y=-2 or y=1
If y=-2 then x=-2+1= -1
If y=1 then x=1+1=2
So, points of intersection are (-1,-2) and (2,1).
Cheers,
Stan H.